OpenStudy (anonymous):

HELP can't find least common multiple of 12,16,24,36

OpenStudy (ashish56):

L.C.M. is 144

OpenStudy (anonymous):

tks how did you find it I tried everything

OpenStudy (saifoo.khan):

Do u want the working?

OpenStudy (anonymous):

yes please

OpenStudy (ashish56):

12\[=\]2\[\times\]2\[\times\]3 16\[=\]2\[\times\]2\[\times\]2\[\times\]2 24\[=\]2\[\times\]2\[\times\]2\[\times\]3 36\[=\]2\[\times\]2\[\times\]3\[\times\]3

OpenStudy (anonymous):

thank you!!!!

OpenStudy (ashish56):

well shown saifoo.khan

OpenStudy (saifoo.khan):

Thank you! ^_^